OpenStudy (anonymous):

whats the distance for (-4,-5)and (1,-2)

OpenStudy (hoblos):

d^2 = (1+4)^2 + (-2+5)^2 = 25 +9 =34 d= sqrt34

OpenStudy (anonymous):

THANKZ